OpenStudy (help_people):

WHOEVER HELPS ME WILLL WIN A MEDAL AND FAN. IT IS A COUPLEL QESTIONS SO....

OpenStudy (anonymous):

I can try to help :)

OpenStudy (acxbox22):

we dont win a fan or medal we earn it :P

OpenStudy (anonymous):

What grade r u in?

OpenStudy (help_people):

The box plots show male and female grades in a U.S. history class: two bar graphs shown. The top one is labeled Male. Minimum at 70, Q1 at 82, median at 88, Q3 at 95. The bottom bar graph is labeled Female. Minimum at 75, Q1 at 78, Q3 at 90, maximum at 92. Which of the following best describes the information about the interquartile ranges? The female interquartile range cannot be known because only three quartiles are showing. The male interquartile range cannot be known because the fourth quartile is missing. The male interquartile range is larger due to the size of the first quartile. The two interquartile ranges are about the same size.

OpenStudy (anonymous):

The interquartile range is quantity Q3 - Q1 We know these values for both males and females. For males: Q3 - Q1 = 95 - 82 = 13 For females: Q3 - Q1 = 90 - 78 = 12

OpenStudy (anonymous):

The two interquartile ranges are about the same size. So, D I believe :)

OpenStudy (help_people):

The box plots show the average daily temperatures in April and October for a U.S. city: two bar graphs shown. The top one is labeled April. Minimum at 50, Q1 at 60, median at 67, Q3 at 71, maximum at 75. The bottom bar graph is labeled October games. Minimum at 45, Q1 at 50, median at 60, Q3 at 75, maximum at 85. What can you tell about the means for these two months? The mean for April is higher than October's mean. There is no way of telling what the means are. The low median for October pulls its mean below April's mean. The high range for October pulls its mean above April's mean.

OpenStudy (anonymous):

I'm not sure! Sorry!

OpenStudy (help_people):

awww well thats to bad @angel12310

OpenStudy (help_people):

@acxbox22 may you help

OpenStudy (anonymous):

I think C

OpenStudy (anonymous):

or D

OpenStudy (help_people):

which one? @angel12310

OpenStudy (help_people):

@acxbox22

OpenStudy (anonymous):

i'm not sure isn't maximum the mean cuz mean means most

OpenStudy (help_people):

how about thus one @angel12310

OpenStudy (help_people):

@acxbox22

OpenStudy (help_people):

The table shows data from a survey about the amount of time students spend doing homework each week. The students were either in college or in high school: High Low Q1 Q3 IQR Median Mean σ College 50 6 8.5 17 8.5 12 15.4 11.7 High School 28 3 4.5 15 10.5 11 10.5 5.8 Which of the choices below best describes how to measure the spread of this data? Both spreads are best described with the IQR. Both spreads are best described with the standard deviation. The college spread is best described by the IQR. The high school spread is best described by the standard deviation. The college spread is best described by the standard deviation. The high school spread is best described by the IQR.